# Fault Diagnosis in New Wind Turbines using Knowledge from Existing Turbines by Generative Domain Adaptation

### v1

@@ Abstract
Intelligent condition monitoring of wind turbines is essential for reducing downtimes. Machine learning models trained on wind turbine operation data are commonly used to detect anomalies and, eventually, operation faults. However, data-driven normal behavior models (NBMs) require a substantial amount of training data, as NBMs trained with scarce data may result in unreliable fault diagnosis. To overcome this limitation, we present a novel generative deep learning approach to make SCADA samples from one wind turbine lacking training data resemble SCADA data from wind turbines with representative training data. Through CycleGAN-based domain mapping, our method enables the application of an NBM trained on an existing wind turbine to one with severely limited data. We demonstrate our approach on field data mapping SCADA samples across 7 substantially different WTs. Our findings show significantly improved fault diagnosis in wind turbines with scarce data. Our method achieves the most similar anomaly scores to an NBM trained with abundant data, outperforming NBMs trained on scarce training data with improvements of +10.3% in F1-score when 1 month of training data is available and +16.8% when 2 weeks are available. The domain mapping approach outperforms conventional fine-tuning at all considered degrees of data scarcity, ranging from 1 to 8 weeks of training data. The proposed technique enables earlier and more reliable fault diagnosis in newly installed wind farms, demonstrating a novel and promising research direction to improve anomaly detection when faced with training data scarcity.
Keywords wind turbine  $\cdot$ condition monitoring  $\cdot$ fault diagnosis  $\cdot$ deep learning  $\cdot$ transfer learning  $\cdot$ generative domain adaptation  $\cdot$ domain mapping
@@ 1 Introduction
Wind energy plays an essential role in the global shift towards renewable energy. With the increasing role of wind energy, improving the reliability of wind turbines (WTs) is of crucial importance [faulstichWindTurbineDowntime2011]. The resulting supervisory control and data acquisition (SCADA) data has facilitated the application of deep learning methods for data-driven condition monitoring tasks in WTs [helbingDeepLearningFault2018a, tautz-weinertUsingSCADAData2017]. Classification models for WT fault diagnosis represent a major share of these models [stetcoMachineLearningMethods2019]. However, these supervised classification methods require thoroughly labeled WT maintenance records and a fault database containing SCADA data for a wide variety of fault types, which is often unavailable in practice. Normal behavior models (NBMs) [helbingDeepLearningFault2018a, meyerMultitargetNormalBehaviour2021] can overcome this limitation. NBMs are based on anomaly detection and model the fault-free operation behavior of WT components. As opposed to classification models, NBMs only require a training dataset comprising historical, fault-free operation measurements. Deviations from the expected value obtained from the NBM at test time, i.e., when the trained model is used in an operational context, can be viewed as an anomaly score, with continuous deviations indicating possible incipient faults.
@@
A major limitation of NBMs is that they require a large training set that is representative of the various WT operation and environmental conditions. Training NBMs with scarce or non-representative data (for instance, only records at low wind speeds) may lead to unreliable fault diagnosis models [grataloupWindTurbineCondition2024, jenkelPrivacyPreservingFleetWideLearning2023]. Large training sets are sometimes unavailable as scarcity of representative training data can occur, for instance, in newly installed WTs, after maintenance, part replacements, or as a result of aging, rendering previously collected data unrepresentative of the future WT operational states. Data scarcity is further exacerbated by the lack of data sharing in the wind industry, which is inhibiting research and development [kusiakRenewablesShareData2016]. A scarcity in training data poses a serious challenge for WT fault diagnosis tasks.
@@
Transfer learning and, in particular, domain adaptation, aims to overcome the limitations caused by data scarcity by transferring knowledge and adapting models from data of a different but related domain [panSurveyTransferLearning2010, wilsonSurveyUnsupervisedDeep2020]. In our fault diagnosis context, domains are typically represented by a WT and its associated data, knowledge, and NBMs. The goal is to use a source domain where data is abundantly available, represented by a WT with abundant data, to improve a task on a target domain, represented by a WT with scarce training data. Domain adaptation has shown success in numerous WT-related fault classification tasks for classifying unlabeled data (e.g., [renNewWindTurbine2019, xieInvestigationDeepTransfer2023, zhuAnomalyDetectionCondition2022]). However, domain adaptation for unsupervised anomaly detection has hardly been investigated [yangAnomalyDetectionDomain2023]. This is likely a result of multiple challenges faced in applying domain adaptation to unsupervised anomaly detection. First, it usually involves time series data as opposed to the emphasis on image data. Further, by definition, only normal data is available for training anomaly detection models, excluding supervision-based methods. Moreover, unlike the typical setting with abundant but unlabeled target domain data, we are faced with data scarcity in the target domain (i.e., the target wind turbine). Available studies using transfer learning or domain adaptation for improving unsupervised anomaly detection through NBMs are limited to fine-tuning [schroderUsingTransferLearning2022] or simple corrections of the NBM [zgraggenTransferLearningApproaches2021].
@@
Performing fault diagnosis for a WT with data scarcity by simply employing an NBM trained on another WT with abundant data is generally not feasible. Since the WTs vary in, for instance, power generation and drive train characteristics, the NBM trained on another WT will predict an incorrect expected state due to the different characteristics learned from its training set. However, if an NBM were employed from a WT exhibiting a minimal domain shift [9238468], i.e., if the 2 WT’s test data distributions would match, it could be used for fault diagnosis for the WT with scarce training data. This would be the case if the two WTs shared the same specifications, operational behavior, and weather conditions. To this end, our work demonstrates how to make the WTs resemble each other by transforming their SCADA data into one another. We employ domain mapping [wilsonSurveyUnsupervisedDeep2020] to map SCADA samples from one WT to resemble the corresponding SCADA data of another WT. Through this mapping, we can enable fault diagnosis using another WT’s NBM. A critical component herein is that the mapping should preserve the sample’s content when translating it to another domain. For example, a SCADA measurement capturing a WT running idly should remain in an idle state, and critically, anomalous behavior should be mapped to anomalous behavior across WT domains.
@@
The authors of [jinConditionMonitoringWind2023] proposed a generative neural network that synthesizes SCADA data resembling another WT but did not employ or consider any content preservation. Without explicitly enforcing a preservation of WT operational states, as employed in our study, anomalies may be mapped to healthy states, ultimately leading to unreliable fault diagnosis. Pattnaik et al. [pattnaikCycleGANBasedUnsupervised2022] presented a content-preserving CycleGAN[zhuUnpairedImagetoImageTranslation2017]-based domain mapping approach for mapping industrial time series data across motors for bearing fault classification. While their work demonstrated the promising potential of mapping data between different but related machines, it followed the usual domain adaptation workflow of abundant but unlabeled data for classification, whereas our study investigates domain mapping for an unsupervised anomaly detection task using limited training data.
@@
We propose a novel fault diagnosis approach based on WT domain mapping with CycleGAN that learns to map SCADA data between the source and the target WT domain in a content-preserving manner. At test time, the target data is mapped to the source domain, thereby enabling fault diagnosis using the source WT’s NBM. To our knowledge, this study is the first to demonstrate a CycleGAN-based framework for unsupervised anomaly detection with time series as well as in wind energy. Our contributions are as follows:
@@
    Our study is the first to investigate and demonstrate domain mapping for unsupervised anomaly detection with time series for WT fault diagnosis.
@@
    We show how our domain mapping technique can be used for data scarce WTs to improve NBM performance in fault diagnosis applications.
@@ 2.1 Domain adaptation
Formally, a domain $D=\{\mathcal{X},P\left(X\right)\}$ consists of a feature space $\mathcal{X}$ and a marginal probability distribution $P(X)$, where $X=\{{x}_{1},\ldots,x_{n}\}\in\mathcal{X}$ [panSurveyTransferLearning2010]. We aim to learn a task $T=\{\mathcal{Y},P(Y|X)\}$, where $\mathcal{Y}$ is the label space and $P(Y|X)$ is the conditional probability. We differentiate between the source domain $\mathcal{D}_{\mathcal{S}}$ with its dataset $D_{s}=\left(x_{1s},y_{1s}\right),\ldots,\left(x_{ns},y_{ns}\right)$ and the target domain $\mathcal{D}_{T}$ with its corresponding dataset $D_{T}=\left(x_{1t},y_{1t}\right),\ldots,\left(x_{nt},y_{nt}\right)$, where usually $0<n_{t}\ll n_{s}$. In our WT fault diagnosis context, the source domain is represented by a source WT with representative SCADA data ($n_{s}\gg 0$) and the target domain by a target WT with scarce, non-representative data ($n_{t}\ll n_{s}$). Transfer learning aspires to use knowledge extracted from $D_{S}$ and by $T_{S}$ to improve learning the task $T_{T}$ in the target domain $\mathcal{D}_{T}\neq\mathcal{D}_{\mathcal{S}}$. Domain adaptation is a particular type of transductive transfer learning, a setting in which the task remains the same but across different domains [panSurveyTransferLearning2010, wilsonSurveyUnsupervisedDeep2020].
@@
As acquiring labels for readily available large scale vision datasets is a prohibitively time intensive and expensive task, much attention has been devoted to unsupervised domain adaptation (labeled source domain and unlabeled target domain data) for vision related classification tasks [wilsonSurveyUnsupervisedDeep2020]. Initial approaches were focused on learning domain-invariant representations of source and target features (e.g., [ganinUnsupervisedDomainAdaptation2015, sunDeepCORALCorrelation2016, tzengDeepDomainConfusion2014]).
@@
More recently, generative methods were able to improve performance in unsupervised domain adaptation [liuDeepUnsupervisedDomain2022, zhaoReviewSingleSourceDeep2022]. These methods extend domain adaptation with a generative component. The main approach is domain mapping (domain translation), which aims to map samples from one domain to another. Domain mapping can allow, e.g., the use of a pretrained source domain classifier for unlabeled target data which has been mapped to the source domain. An essential challenge herein is to retain important characteristics and semantics of the mapping input such as class characteristics, to ensure a correct and consistent classification. For instance, a handwritten digit of class 1 must at the very least remain of class 1 when mapped into a corresponding street view house numbers domain for a correct classification.
Unpaired image-to-image translation methods (e.g., [liuUnsupervisedImagetoImageTranslation2017, zhuUnpairedImagetoImageTranslation2017]) are fundamental to achieve such a content-preserving mapping. CycleGAN [zhuUnpairedImagetoImageTranslation2017] enables the translation of unpaired samples into another domain with generative adversarial networks through a cycle-consistency loss, which enforces the preservation of content. Domain mapping based on CycleGAN has been used in various applications, e.g. in medical settings [palladinoUnsupervisedDomainAdaptation2020], voice conversion [kanekoParallelDataFreeVoiceConversion2017], as well as machine fault diagnosis [pattnaikCycleGANBasedUnsupervised2022]. For a more comprehensive review of discrepancy-based domain adaptation and domain mapping we refer to [liuDeepUnsupervisedDomain2022, zhaoReviewSingleSourceDeep2022].
2.2 Domain adaptation for wind turbine fault diagnosis
@@
The potential of domain adaptation has been demonstrated with time series data from sensors across multiple applications, including non-WT industrial fault detection [pattnaikCycleGANBasedUnsupervised2022, shiDeepUnsupervisedDomain2022, yanComprehensiveSurveyDeep2024]. These applications primarily rely on classification through domain alignment with abundant target domain data lacking fault labels. For WTs specifically, discrepancy-based domain adaptation for fault classification is presented in [renNewWindTurbine2019, xieInvestigationDeepTransfer2023, yueSpatioTemporalFeatureAlignment2024, zhuAnomalyDetectionCondition2022]. Results show significant gains in accuracy on the unlabeled target domains, even when target domain data is scarce [xieInvestigationDeepTransfer2023]. An alternative approach to improve fault classification performance, namely by generative data augmentation, is proposed in [liuNovelTransferLearning2022] to generate missing working conditions for vibration data.
@@
Yet, the available literature for our relevant anomaly detection task using NBMs is scarce. In our unsupervised anomaly detection setting, there is typically a target data scarcity, coupled with no available supervision (e.g., no fault labels), thereby rendering many previously presented domain adaptation techniques unsuitable for this task. Few conventional transfer learning approaches have been proposed: In [schroderUsingTransferLearning2022], fine-tuning is successfully applied to an NBM pretrained on physics-informed simulation data. While the results show significantly improved performance for when only one month of SCADA data is available, the physics-informed simulation restricts the models to active power NBMs. A simple correction based on linear regression of the source domain NBM output is proposed in [zgraggenTransferLearningApproaches2021]. The proposed correction achieves slightly improved results over fine-tuning but is primarily implemented to avoid overfitting an NBM on a specific season, represented by a target training set comprising considerable 3 months of data.
@@
The first generative domain adaptation work for WT NBMs is presented by Jin et al. [jinConditionMonitoringWind2023]. Specifically, a generative adversarial network (GAN) is proposed to map samples from the scarce target domain, consisting of two weeks of SCADA data, to the source domain comprising data from a different WT. The GAN is conditioned on target data, that is, the input to the generator are SCADA samples from the target domain, rather than random noise. In doing so, the target domain input is mapped to samples matching the marginal probability distribution of the source domain. Finally, the mapped target data is used with an autoencoder-based NBM pretrained on the source domain for anomaly detection. Presented results demonstrate reliable anomaly detection despite target data scarcity. While this approach resembles domain mapping, it substantially deviates from previously outlined approaches. First, it consists of only one-directional target-to-source mappings. Crucially, the mapping network is also unconstrained, i.e., there is no constraint to preserve specific WT states (e.g., maximal power generation or anomalous operation behavior) to the corresponding state in the other domain. This may result in anomalous data being mapped to arbitrary healthy operation states. This lack of content preservation can therefore lead to unreliable fault diagnosis. Our study overcomes such deficiencies by employing domain mapping based on CycleGAN with consistency losses to preserve the SCADA content during domain translation.
As the only study to investigate constrained domain mapping for WTs, Chatterjee et al. [chatterjeeDomaininvariantIcingDetection2023] show only a marginal and negligible accuracy gain in blade icing detection when using CycleGAN to map images of plain ice-free rotor blades across a dataset of rotor blades with icing for generative data augmentation. The authors attributed the poor performance to possibly the small dataset size, large class imbalance, or their image generation process. Their work utilized images of rotor blades with all modes present and for synthetic data augmentation (i.e., convert images of plain blades to images of blades affected by icing), whereas our study investigates the translation of non-representative SCADA time series data for anomaly detection.

### v2

# Fault Detection in New Wind Turbines with Limited Data by Generative Transfer Learning

@@ Abstract
Intelligent condition monitoring of wind turbines is essential for reducing downtimes. Machine learning models trained on wind turbine operation data are commonly used to detect anomalies and, eventually, operation faults. However, data-driven normal behavior models (NBMs) require a substantial amount of training data, as NBMs trained with scarce data may result in unreliable fault detection. To overcome this limitation, we present a novel generative deep transfer learning approach to make SCADA samples from one wind turbine lacking training data resemble SCADA data from wind turbines with representative training data. Through CycleGAN-based domain mapping, our method enables the application of an NBM trained on an existing wind turbine to a new one with severely limited data. We demonstrate our approach on field data mapping SCADA samples across 7 substantially different WTs. Our findings show significantly improved fault detection in wind turbines with scarce data. Our method achieves the most similar anomaly scores to an NBM trained with abundant data, outperforming NBMs trained on scarce training data with improvements of +10.3% in F1-score when 1 month of training data is available and +16.8% when 2 weeks are available. The domain mapping approach outperforms conventional fine-tuning at all considered degrees of data scarcity, ranging from 1 to 8 weeks of training data. The proposed technique enables earlier and more reliable fault detection in newly installed wind farms, demonstrating a novel and promising research direction to improve anomaly detection when faced with training data scarcity.
Keywords wind turbine  $\cdot$ fault detection  $\cdot$ deep learning  $\cdot$ transfer learning  $\cdot$ generative domain adaptation  $\cdot$ domain mapping  $\cdot$ data scarcity  $\cdot$ anomaly detection
@@ 1 Introduction
Wind energy plays an essential role in the global shift towards renewable energy. With the increasing role of wind energy, improving the reliability of wind turbines (WTs) is of crucial importance [1]. Supervisory control and data acquisition (SCADA) data has facilitated the application of deep learning methods for data-driven condition monitoring tasks in WTs [2, 3]. Classification models represent a major share of these models [4]. However, these supervised classification methods require thoroughly labeled WT maintenance records and a fault database containing SCADA data for a wide variety of fault types, which is often unavailable in practice. Normal behavior models (NBMs) [2, 5] can overcome this limitation. NBMs are based on anomaly detection and model the fault-free operation behavior of WT components. As opposed to classification models, NBMs only require a training dataset comprising historical, fault-free operation measurements. Deviations from the expected value obtained from the NBM at test time, i.e., when the trained model is used in an operational context, can be viewed as an anomaly score, with continuous deviations indicating possible incipient faults. This setting of training data from only assumed normal operation falls under unsupervised anomaly detection, a category that is in literature also sometimes referred to as semi-supervised [6].
@@
A major limitation of NBMs is that they require a large training set that is representative of the various WT operation and environmental conditions. Training NBMs with scarce or non-representative data (for instance, only records at low wind speeds) may lead to unreliable fault detection models [7, 8]. Large training sets are sometimes unavailable as scarcity of representative training data can occur, for instance, in newly installed WTs, after maintenance, part replacements, or as a result of aging, rendering previously collected data unrepresentative of the future WT operational states. Data scarcity is further exacerbated by the lack of data sharing in the wind industry, which is inhibiting research and development [9]. A scarcity in training data poses a serious challenge for WT fault detection tasks.
@@
Transfer learning and, in particular, domain adaptation, aims to overcome the limitations caused by data scarcity by transferring knowledge and adapting models from data of a different but related domain [10, 11]. In our fault detection context, domains are represented by a WT and its associated data, knowledge, and NBMs. The goal is to use a source domain where data is abundantly available, represented by a WT with abundant data, to improve a task on a target domain, represented by a WT with scarce training data. Domain adaptation has shown success in numerous WT-related fault classification tasks for classifying unlabeled data (e.g., [12, 13, 14]). However, domain adaptation for unsupervised anomaly detection has hardly been investigated [15]. This is likely a result of multiple challenges faced in applying domain adaptation to unsupervised anomaly detection. First, it usually involves time series data as opposed to the emphasis on image data. Further, by definition, only normal data is available for training anomaly detection models, excluding supervision-based methods. Moreover, unlike the typical setting with abundant but unlabeled target domain data, we are faced with data scarcity in the target domain (i.e., the target wind turbine). Available studies using transfer learning or domain adaptation for improving unsupervised anomaly detection through NBMs are limited to fine-tuning [16] or simple corrections of the NBM [17].
@@
Adapting and transferring knowledge across WTs is essential. Detecting faults in a data-scarce WT by employing an NBM trained on another data-rich WT is generally not feasible. Since the WTs vary in, for instance, power generation and drive train characteristics, the NBM trained on a different WT will predict an incorrect expected state due to the varying characteristics. However, an NBM from another WT exhibiting a minimal domain shift [18] might possibly be used to detect faults in the WT with scarce training data. That is, if the 2 WT’s test data distributions would match. This would be the case if the two WTs shared the same specifications, operational behavior, and weather conditions. To this end, our work demonstrates how to make the WTs resemble each other by transforming their SCADA data into one another. We employ domain mapping [11] to map SCADA samples from one WT to resemble the corresponding SCADA data of another WT. Through this mapping, we can enable fault detection using another WT’s NBM. A critical component herein is that the mapping should preserve the sample’s content when translating it to another domain. For example, a SCADA measurement capturing a WT running idly should remain in an idle state, and critically, anomalous behavior should be mapped to anomalous behavior across WT domains.
@@
Jin et al. [19] proposed a generative neural network that synthesizes SCADA data resembling another WT but did not employ or consider any content preservation. Without explicitly enforcing a preservation of WT operational states, as employed in our study, anomalies may be mapped to healthy states, ultimately leading to unreliable fault detection. Pattnaik et al. [20] presented a content-preserving CycleGAN [21]-based domain mapping approach for mapping industrial time series data across motors for bearing fault classification. While their work demonstrated the promising potential of mapping data between different but related machines, it followed the usual domain adaptation workflow of abundant but unlabeled data for classification, whereas our study investigates domain mapping for an unsupervised anomaly detection task using limited training data.
@@
We propose a novel fault detection approach based on WT domain mapping with CycleGAN that learns to map SCADA data between the source and the target WT domain in a content-preserving manner. At test time, the target data is mapped to the source domain, thereby enabling fault detection using the source WT’s NBM. To our knowledge, this study is the first to demonstrate a CycleGAN-based framework for unsupervised anomaly detection with time series as well as in wind energy. Our contributions are as follows:
@@
    Our study is the first to investigate and demonstrate domain mapping for unsupervised anomaly detection with time series for WT fault detection.
@@
    We show how our domain mapping technique can be used for data scarce WTs to improve NBM performance in fault detection applications.
@@ 2.1 Domain adaptation
A promising approach to overcome data scarcity limitations is transfer learning, particularly domain adaptation [10, 11]. The goal is adapt data, models, or knowledge from one domain with abundant data, the source domain, to a domain affected by some data or label scarcity, the target domain.
@@
Formally, a domain $D=\{\mathcal{X},P\left(X\right)\}$ consists of a feature space $\mathcal{X}$ and a marginal probability distribution $P(X)$, where $X=\{{x}_{1},\ldots,x_{n}\}\in\mathcal{X}$ [10]. We aim to learn a task $T=\{\mathcal{Y},P(Y|X)\}$, where $\mathcal{Y}$ is the label space and $P(Y|X)$ is the conditional probability. We differentiate between the source domain $\mathcal{D}_{\mathcal{S}}$ with its dataset $D_{s}=\left(x_{1s},y_{1s}\right),\ldots,\left(x_{ns},y_{ns}\right)$ and the target domain $\mathcal{D}_{T}$ with its corresponding dataset $D_{T}=\left(x_{1t},y_{1t}\right),\ldots,\left(x_{nt},y_{nt}\right)$, where usually $0<n_{t}\ll n_{s}$. In our context, the source domain is represented by a source WT with representative SCADA data ($n_{s}\gg 0$) and the target domain by a target WT with scarce data ($n_{t}\ll n_{s}$). Transfer learning aspires to use knowledge extracted from $D_{S}$ and by $T_{S}$ to improve learning $T_{T}$ in the target domain $\mathcal{D}_{T}\neq\mathcal{D}_{\mathcal{S}}$. Domain adaptation is a particular type of transductive transfer learning, a setting in which the task remains the same but across different domains [10, 11].
@@
Much attention has been devoted to unsupervised domain adaptation, a setting with a labeled source domain but an unlabeled target domain, since this is common for vision datasets where acquiring labels is a time intensive and expensive task [11]. Initial approaches were focused on learning domain-invariant representations of source and target features to enable classification of the unlabeled target domain (e.g., [22, 23]).
@@
More recently, generative methods were able to improve performance in unsupervised domain adaptation [24, 18]. These methods extend domain adaptation with a generative component. A main approach is domain mapping, which aims to map (translate) samples from one domain to another. Domain mapping can allow the use of a source domain classifier for unlabeled mapped target data. As an example, we consider datasets of digits. Given unlabeled handwritten digits, no classifier can be trained. But with domain mapping, they can be mapped even without defined pairs to resemble another domain, e.g., computer-generated numbers, for which labels and a classifier are available. An essential challenge is to retain important characteristics of the input, such as its class, to ensure a correct classification. That is, a handwritten digit 7 must remain a 7 when mapped. The underlying unpaired image-to-image translation methods (e.g., [25, 21]) are fundamental to achieve this. CycleGAN [21] enables the translation of unpaired samples into another domain with generative adversarial networks through a cycle-consistency loss enforcing the preservation of content. Domain mapping with CycleGAN has been used in various applications, e.g. in medical settings [26], voice conversion [27], machine fault diagnosis [20], or also for synthetic data augmentation in WT icing detection [28]. For a more comprehensive review of unsupervised domain adaptation we refer to [24, 18].
2.2 Domain adaptation for wind turbine fault detection
@@
The potential of domain adaptation has been demonstrated with time series data from sensors across multiple applications, including non-WT industrial fault detection [20, 29, 30]. These applications primarily rely on classification through domain alignment with abundant target domain data lacking fault labels. For WTs specifically, discrepancy-based domain adaptation for fault classification is presented in [12, 13, 31, 14]. Results show significant gains in accuracy on the unlabeled target domains, even when target domain data is scarce [13]. An alternative approach to improve fault classification performance, namely by generative data augmentation, is proposed in [32] to generate missing working conditions for vibration data.
@@
Yet, the available literature for our relevant anomaly detection task using NBMs is scarce. In our unsupervised anomaly detection setting, there is typically a target data scarcity, coupled with no available supervision (e.g., no fault labels), thereby rendering many previously presented domain adaptation techniques unsuitable for this task. Few conventional transfer learning approaches have been proposed: In [16], fine-tuning is successfully applied to an NBM pretrained on physics-informed simulation data. While the results show significantly improved performance for when only one month of SCADA data is available, the physics-informed simulation restricts the models to active power NBMs. A simple correction based on linear regression of the source domain NBM output is proposed in [17]. The proposed correction achieves slightly improved results over fine-tuning but is primarily implemented to avoid overfitting an NBM on a specific season, represented by a target training set comprising considerable 3 months of data.
@@
The first generative domain adaptation work for WT NBMs is presented by Jin et al. [19]. Specifically, a generative adversarial network (GAN) is proposed to map samples from a data-scarce WT, comprising two weeks of SCADA data, to a data-rich WT. Like in domain mapping approaches, the GAN is conditioned on target data, that is, the input to the generator are SCADA samples from the scarce target domain, rather than random noise. In doing so, the target domain input is mapped to samples matching the marginal probability distribution of the source domain. Finally, the mapped target data is used with an autoencoder-based NBM pretrained on the source domain for anomaly detection. Presented results demonstrate more reliable anomaly detection despite target data scarcity. While this approach resembles domain mapping, it substantially deviates from its approaches. For one, it consists of only one-directional target-to-source mappings. Crucially, the mapping network is also unconstrained, i.e., there is no constraint to preserve specific WT states (e.g., maximal power generation or anomalous behavior) to their corresponding state in the other domain. This may result in anomalous data being mapped to arbitrary healthy operation states. This lack of content preservation could therefore lead to unreliable fault detection. Our study overcomes such deficiencies by employing domain mapping based on CycleGAN with consistency losses to preserve the SCADA content during domain translation. Additionally, our work incorporates multi-hour long SCADA time series, as opposed to features from only a single time point.